3 The Stark Company manufactures a product at s expected to our 20 per unit n variable roduct on costs and se or 4) er unit. The sales o m sso s 0% o the sales n ue mense competition, Stark actua so o units for $38 per unit. The actual variable production costs incurred were $23.75 per unit. Calculate the total contribution margin and contribution margin ratio at the expected price/costs and the actual price/costs. How might management use this formation? Complete the expected contribution margin by entering the appropriate amounts to calculate the total contribution margin and contribution margin ratio. (Enter the contribution margin ratio to two decimal place, XXX%.) Expected Contribution Margin Net Sales Revenue Variable Costs: Production Costs Sales Commission Contribution Margin Contribution Margin Ratio Complete the actual contribution margin by entering the appropriate amounts to calculate the total contribution margin and contribution margin ratio. (Enter the contribution margin ratio to two decimal place, Xxx%.) Actual Contribution Margin Net Sales Revenue Variable Costs: Production Costs Sales Commission Contribution Margin Contribution Margin Ratio
Answer #1

Expected Contribution Margin Net sales revenue (200 S40) Variable cost: Production costs 200 $20 Sales Commission 200 * $4 Contribution margin Contribution margin Ratio (3,200/8000)* 100 4,000 800 4,800 3,200 40.00% Actual Contribution Margin Net sales revenue (200 38 7,600 Variable cost: Production costs 200 $23.75 Sales Commission 200 $3.8 Contribution margin Contribution margin Ratio (2090/7600) 100 4,750 760 5,510 2,090 27.50%

Expected Sales commission = 40 * 10 / 100 = $4

Actual sales commission = 38 * 10 / 100 = $3.8

Contribution = Net sales revenue - (variable production cost + Sales Commission)

Due to competition the company reduced the sales price and at the same time variable cost is increased so the company contribution margin was reduced as compared to actual with the Expected. So the Actual net income of the company also will be low as compared to the Expected Net Income.